The following formulas represent the demand and supply curves for corn:
QD = 1,600 – 125 * P
QS = 440 + 165 * P
Calculate the equilibrium price and quantity in this market and illustrate this graphically. Suppose corn becomes less popular so the market demand curve is now given by QD = 1,020 – 125 * P. Calculate the new equilibrium price and quantity and illustrate the movement from the old equilibrium to the new one graphically.
Answer #1

Answer
the equilibrium is at Qd=Qs
1600-125P=440+165P
290P=1160
P=4
Q=1600-125*4
Q=1100
the equilibrium price is $4, and the quantity is 1100 units

after the change in demand
1020-125P=440+165P
290P=580
P=2
Q=1020-125*2
Q=770
the new price is $2, and the quantity is 770 units

price (0, 12.8) 10 (0, 8.16) demand old su old equilibrium demand new (1100, 4) new equilibriu (770, 2) (1020, 0) (1600, 0) 2